# Charged spherically symmetric and slowly rotating charged black hole solutions in bumblebee gravity

### v1

@@ IV.2 Case B: $V(X)=\frac{\lambda}{2}X$ and $\Lambda\neq 0$
By calculating the errors in the field equations, we find that the results are the same as case A. Similarly, the rotating solution must satisfy the condition that $a^{2}\ell$ is a small quantity.
@@ V SHADOW OBSERVABLES
The black hole shadow is formed by the strong gravitational field of the black hole. The shadow appears as a two-dimensional dark region to a distant observer. Shadows serve as one of the most intuitive pieces of evidence for the presence of black holes. Particularly, the unveiling of an image provides even stronger evidence for black hole’s existence Akiyama et al. (2019a, b, c, d); Akiyama et al. (2022). The shadow of the Kerr-like black hole in bumblebee gravity has been extensively studied in previous works Wang and Wei (2022); Ding et al. (2020); Kuang and Övgün (2022); Chen et al. (2020). In this section, we aim to investigate the shadow of the KR-AdS like black holes in the framework of Einstein-bumblebee gravity.
@@ V.1 The photon orbit
In this section, we will derive the equations of motion of photons orbiting the black holes. From Eqs. (61) and (62), it is apparent that the field equations hold when the rotating black hole solution satisfies that $a\ell^{2}$ is a small quantity. Therefore, in the subsequent part, all the analysies are under this condition. The motion of a photon in the background of the black hole is given by the geodesics. However, the process of solving geodesic equations is very difficult. we choose to use the Hamilton-Jacobi approach to solve it. The Lagrangian for the photon is
@@
$V_{eff}=-\frac{E^{2}}{(1+\ell)}\left[\left[\left(r^{2}+(1+\ell)a^{2}\right)-\sqrt{1+\ell}a\xi\Xi\right]^{2}-\Delta_{r}\big{[}\frac{\left(\sqrt{1+\ell}a-\xi\Xi\right)^{2}}{\Delta_{\theta}}+\eta\big{]}\right].$ (77)
In Fig. 1, we depict the effective potential $V_{\text{eff}}$ for a rotating charged black hole with $\ell=0$ and $\ell=0.1$. From Fig. 1, we can observe that Lorentz-violating parameter greater than 0 can increase the effective potential Wei et al. (2015). For the unstable circular null orbit, the effective potential satisfies
@@
Then we can obtain
@@ V.2 Observables
$x$ $=$ $\lim_{r\to\infty}\frac{rp^{(\theta)}}{p^{(t)}}=\sqrt{\eta+a^{2}\cos^{2}\theta-\xi^{2}\cot^{2}\theta},$ (81)
$y$ $=$ $\lim_{r\to\infty}\frac{-rp^{(\varphi)}}{p^{(t)}}=-\xi\csc\theta.$ (82)
@@
It is easy to see that both the coordinates $x$ and $y$ are the functions of $(M,Q_{eff},a,\Lambda_{e},\ell,\theta)$. We set $M=1$, and show the shapes of the shadow in Fig. 2. Similar to the Kerr-Newman-AdS black hole, the shadow radius increases with the negative effective cosmological costant $\Lambda_{e}$, as shown in Fig. 2(a). From Fig. 2(b) we can see that, the position of the shadow depends on the observation angle $\theta$. The Lorentz-violating parameter $\ell$ also affects the shaow size which can be seen from Fig. 2(c). Besides, the deformation of the shadow becomes more pronounced when the charge parameter $Q_{0}$ increases, which can be seen from Fig. 2(d).
@@ VI CONCLUSIONS
Moreover, in the background of the KN-(A)dS-like black hole, we derived null geodesics using the Hamilton-Jacobi equation. Utilizing the conditions for unstable circular null orbits, we derived two conserved parameters, $\xi$ and $\eta$, which govern the photon’s motion. Using these parameters, we determined the celestial coordinates of the shadow. Subsequently, we investigated the size and deformation of the shadow. By varying the observation angle $\theta=\frac{\pi}{2},$ $\theta=\frac{\pi}{4},$ $\theta=\frac{\pi}{6}$, we calculated the shadow’s radius $R_{s}$ for different parameters. We observed that for fixed spin $a$ and Lorentz-violating parameter $\ell$, observers situated at lower latitudes consistently observe a larger shadow. Meanwhile, we found that the radius of the reference circle decreases with the Lorentz-violating parameter and electric charge , while the distortion parameter increases with the Lorentz-violating parameter and electric charge.
@@ Acknowledgements
We are grateful to Tao-tao Sui and Ke Yang for useful discussions. This work was supported by the National Key Research and Development Program of China (Grant No. 2021YFC2203003), the National Natural Science Foundation of China (Grants No. 12205129, No. 12147166, and No. 12247101), the China Postdoctoral Science Foundation (Grant No. 2021M701529), the 111 Project (Grant No. B20063), the Major Science and Technology Projects of Gansu Province, and Lanzhou City’s scientific research funding subsidy to Lanzhou University.

### v2

@@ IV.2 Case B: $V(X)=\frac{\lambda}{2}X$ and $\Lambda\neq 0$
By calculating the errors in the field equations, we find that the results are similar to case A. Similarly, the rotating solution must satisfy the condition that $a^{2}\ell$ is a small quantity.
@@ V SHADOW OBSERVABLES
The black hole shadow is formed by the strong gravitational field of the black hole. The shadow appears as a two-dimensional dark region to a distant observer. Shadows serve as one of the most intuitive pieces of evidence for the presence of black holes. Particularly, the unveiling of an image provides even stronger evidence for black hole’s existence Akiyama et al. (2019a, b, c, d); Akiyama et al. (2022). The shadow of the Kerr-like black hole in bumblebee gravity has been extensively studied in previous works Wang and Wei (2022); Ding et al. (2020); Kuang and Övgün (2022); Chen et al. (2020); Islam et al. (2024); Afrin et al. (2024). In this section, we aim to investigate the shadow of the KN-AdS like black holes in the framework of Einstein-bumblebee gravity.
@@ V.1 The photon orbit
In this section, we will derive the equations of motion of photons orbiting the black holes. From Eqs. (63) and (64), it is apparent that the field equations hold when the rotating black hole solution satisfies that $a\ell^{2}$ is a small quantity. Therefore, in the subsequent part, all the analysies are under this condition. The motion of a photon in the background of the black hole is given by the geodesics. However, the process of solving geodesic equations is very difficult, we choose to use the Hamilton-Jacobi approach to solve it. The Lagrangian for the photon is
@@
$V_{\text{eff}}=-\frac{E^{2}}{(1+\ell)}\left[\left[\left(r^{2}+(1+\ell)a^{2}\right)-\sqrt{1+\ell}a\xi\Xi\right]^{2}-\Delta_{r}\big{[}\frac{\left(\sqrt{1+\ell}a-\xi\Xi\right)^{2}}{\Delta_{\theta}}+\eta\big{]}\right].$ (79)
In Fig. 3, we depict the effective potential $V_{\text{eff}}$ for a rotating charged black hole with $\ell=0$ and $\ell=0.1$. From Fig. 3, we can observe that Lorentz-violating parameter greater than 0 can increase the effective potential Wei et al. (2015b). For the unstable circular null orbit, the effective potential satisfies
@@
then we can obtain Tsukamoto et al. (2014); Tsukamoto (2018)
@@ V.2 Observables
$x$ $=$ $\lim_{r\to\infty}\frac{-rp^{(\varphi)}}{p^{(t)}}=-\xi\csc\theta,$ (83)
$y$ $=$ $\lim_{r\to\infty}\frac{rp^{(\theta)}}{p^{(t)}}=\pm\sqrt{\eta+a^{2}\cos^{2}\theta-\xi^{2}\cot^{2}\theta}.$ (84)
@@
It is easy to see that both the coordinates $x$ and $y$ are the functions of $(M,Q_{eff},a,\Lambda_{e},\ell,\theta)$. We set $M=1$, and show the shapes of the shadow in Fig. 4. Similar to the KN-AdS black hole, the shadow radius increases with the negative effective cosmological costant $\Lambda_{e}$, as shown in Fig. 4(a). From Fig. 4(b) we can see that, the position of the shadow depends on the observation angle $\theta$. The Lorentz-violating parameter $\ell$ also affects the shaow size which can be seen from Fig. 4(c). Besides, the deformation of the shadow becomes more pronounced when the charge parameter $Q_{0}$ increases, which can be seen from Fig. 4(d).
@@
The shadow image of the supermassive black hole M87* was photographed by the EHT collaboration with the mass $M=6.5\times 10^{9}M_{\odot}$, the distance $r_{0}=16.8Mpc$, and the inclination angle $\theta_{0}=17^{o}$ Akiyama et al. (2019a, b, c, d); Akiyama et al. (2022). And in Ref. Kocherlakota et al. (2021), the authors stated that the shadow size of M87* should lie between $4.31M$ and $6.08M$. We will use the shadow observable $R_{a}$ to constrain the parameters of the Kerr-Newman-like black hole in bumblebee gravity. The definition of the characteristic areal radius of the shadow curve can be expressed as Banerjee et al. (2020); Abdujabbarov et al. (2015)
$R_{a}=\sqrt{\frac{2}{\pi}\int^{r_{max}}_{r_{min}}\Big{(}y(r)\frac{dx(r)}{dr}\Big{)}}.$ (88)
Figure 9: The constraints of the EHT observation $4.31M\leq R_{a}\leq 6.08M$ on the parameters ($a,Q_{0}$). The real curves represent $R_{a}=4.31M$ with different Lorentz-violating parameters $\ell$.
Figure 9 shows the constraints on the black hole parameters by the EHT observation with the shadow size $R_{a}$. All the solid lines represent the lower bound $R_{a}=4.31M$ or $\frac{(1+\ell)(2Q_{0}^{2}+(2+\ell)a^{2})}{2+\ell}=M^{2}$. The reasonable parameters, subject to the constraints, lie below these lines. The results illustrate that the larger the Lorentz-violating parameter $\ell$, the stronger the constraint on the rotating parameter $a$ and charge parameter $Q_{0}$. For $\ell=-0.1$ and $\ell=0.1$, the ranges of the charge paramter $Q_{0}$ are $Q_{0}/M\in(0,0.93)$ and $Q_{0}/M\in(0,0.88)$, respectively. When we consider a small Lorentz-violating parameter, according to the range of the shadow size $R_{a}$, we can not make any constraint on the rotating parameter $a$.
@@ VI CONCLUSIONS
Moreover, in the background of the KN-(A)dS-like black hole, we derived null geodesics using the Hamilton-Jacobi equation. Utilizing the conditions for unstable circular null orbits, we derived two conserved parameters, $\xi$ and $\eta$, which govern the photon’s motion. Using these parameters, we determined the celestial coordinates of the shadow. Subsequently, we investigated the size and deformation of the shadow. By varying the observation angle $\theta=\frac{\pi}{2},$ $\theta=\frac{\pi}{4},$ $\theta=\frac{\pi}{6}$, we calculated the shadow’s radius $R_{s}$ for different parameters. We observed that for fixed spin parameter $a$ and the Lorentz-violating parameter $\ell$, observers situated at lower latitudes consistently observe a larger shadow. Meanwhile, we found that the radius of the reference circle decreases with the Lorentz-violating parameter and electric charge , while the distortion parameter increases with the Lorentz-violating parameter and electric charge.
@@ Acknowledgements
We are grateful to Tao-Tao Sui and Ke Yang for useful discussions. This work was supported by the National Key Research and Development Program of China (Grant No. 2021YFC2203003), the National Natural Science Foundation of China (Grants No. 12475056, No. 12247101, No. 12475055 and No. 12205129), the 111 Project (Grant No. B20063), Gansu Province’s Top Leading Talent Support Plan.
Appendix A A BRIEF DISCUSSION ON BLACK HOLE THERMODYNAMICS
In this appendix, we briefly discuss the impact of $\ell$ $(\ell=\xi b^{2})$ on black hole thermodynamics. First, using the surface gravity, the black hole temperature is given by
$T_{h}=\frac{-2{Q_{0}}^{2}(1+\ell)-{r_{h}}^{2}(2+\ell)\left(\Lambda_{e}{r_{h}}^{2}(1+\ell)-1\right)}{4\pi{r_{h}}^{3}\sqrt{1+\ell}(2+\ell)},$ (89)
where $r_{h}$ is the event horizon radius. Noting that when $\left|\ell\right|<1$, we can deduce that by redefining certain physical quantities ($Q=\sqrt{\frac{2(1+\ell)}{2+\ell}}{Q_{0}}$, $\Lambda=(1+\ell)\Lambda_{e}$, $T_{h}^{\prime}=\sqrt{1+\ell}T_{h}$), the black hole temperature can be restored to the original form of the RN-AdS black hole. This means that the curve of the temperature as a function of the horizon radius does not change fundamentally. For the RN-AdS-like black hole in bumblebee gravity, the extended first law of thermodynamics can be formulated as
$dM=T_{h}dS+\Phi_{0}dQ_{0}+\frac{\Theta}{8\pi}d\Lambda_{e},$ (90)
where $S=4\pi\sqrt{1+\ell}r_{h}^{2}$, $\Theta=-\frac{4}{3}(1+\ell)\pi r_{h}^{3}$, $\Phi_{0}=\frac{2(1+\ell)Q_{0}}{(2+\ell)r}$. This is similar to the first law of thermodynamics for the RN-AdS black hole.
Furthermore, by using the critical point equations
$\partial_{r_{h}}T\left(r_{h}\right)=0,\quad\partial_{r_{h}}^{2}T\left(r_{h}\right)=0,$ (91)
we can easily obtain $Q_{0}=\frac{\sqrt{2+\ell}r_{h}}{\sqrt{6(1+\ell)}}$ and $\Lambda_{e}=\frac{1}{2{r_{h}}^{2}(1+\ell)}$ which only differ from the critical values in the RN-AdS black hole by a coefficient. Therefore, the critical point has indeed been modified, but there is no essential difference from the RN-AdS black hole.
The specific heat can be calculated as
$C_{\text{B}}=T_{h}\left(\frac{\partial S}{\partial T_{h}}\right)_{Q_{0},\Lambda_{e}}=\frac{2\pi\sqrt{1+\ell}r_{h}^{2}\left(-\left(\Lambda_{e}(1+\ell)r_{h}^{4}\right)+r_{h}^{2}-\frac{2{Q_{0}}^{2}(1+\ell)}{\ell+2}\right)}{\frac{6{Q_{0}}^{2}(1+\ell)}{2+\ell}-r_{h}^{2}\left(\Lambda_{e}(1+\ell)r_{h}^{2}+1\right)}.$ (92)
It can be seen that the specific heat has been modified. By redefining certain physical quantities ($Q=\sqrt{\frac{2(1+\ell)}{2+\ell}}{Q_{0}}$, $\Lambda=(1+\ell)\Lambda_{e}$), the specific heat becomes equal to that of the RN-AdS black hole, multiplied by a factor, satisfying $C_{\text{B}}=\sqrt{1+\ell}C_{\text{R}}$. It is found that for $01+12Y>$ 0, where $Y=\frac{(1+\ell)^{2}{Q_{0}}^{2}\Lambda_{e}}{(1+\frac{\ell}{2})}$, the heat capacity is positive within the intervals $r_{h}\subset\left(\sqrt{\frac{\sqrt{1-4Y}-1}{-2(1+\ell)\Lambda_{e}}},\sqrt{\frac{1-\sqrt{1+12Y}}{-2(1+\ell)\Lambda_{e}}}\right)\cup\left(\sqrt{\frac{1+\sqrt{1+12Y}}{-2(1+\ell)\Lambda_{e}}},\infty\right)$. While $1+12Y\leq 0$, the heat capacity is positive for $r_{h}\subset\left(\sqrt{\frac{\sqrt{1-4Y}-1}{-2(1+\ell)\Lambda_{e}}},\infty\right)$. From this, we conclude that stability of black hole thermodynamics does not change fundamentally. On the other hand, from the black hole free energy
$F=M-T_{h}S=\frac{1}{4}\left(\frac{3(1+\ell)Q_{0}^{2}}{\left(1+\frac{\ell}{2}\right)r_{h}}+3(1+\ell)\Lambda_{e}r_{h}^{3}+r_{h}\right),$ (93)
it reveals the possible black hole phase transitions are similar to those of the RN-AdS black hole, the system may exhibit black hole phases with the following stability: small black holes (stable), medium black holes (unstable), and large black holes (stable).
From these aspects, we can conclude that the nonminimal coupling does indeed modify the black hole thermodynamics, but it does not change some of the essential behaviors. Additionally, the modification parameter is generally small, indicating that Lorentz symmetry breaking has a relatively minor impact on black hole thermodynamics in this model.